Edwards Construction currently has debt outstanding with a market value of $92,000 and a cost of 7 percent. The company has Ea-1. a-2. Value of equity Debt-to-value ratio Equity value Debt-to-value Equity value Debt-to-value
Answer #1

Ans:-(a-1) In this part we need to find the value of the equity. For that first, we need to find the value of the firm

The market value of outstanding debt is given $92,000 and the cost is given 7%.

Now the value of the firm will be given by EBIT / Cost = $6440 / 7% = $92,000.

Therefore the value of the firm is equal to $92,000.

Now the value of equity = Value of firm - Value of debt = $92,000 - $92,000 = 0.

Therefore the value of equity is 0.

(a-2) Debt to value ratio is given by Debt / Value of the firm = $92,000 / $92,000 = 1.

(c) In this part, we need to find the equity value and debt to value ratio For that first, we need to find the value of equity

If the growth rate is 2% then the Earnings next year will be EBIT * ( 1 + g), where g is the growth rate.

= $6,440 * ( 1 + 0.02) = $6568.80.

Now the payment to shareholders will be (Earnings next year - EBIT = $6568.80 - $6,440 = $128.80)

Now the value of equity will be Payment to shareholders / ( Rate of return - Growth rate)

= $128.80 / ( 0.07 - 0.02) = $2576.

Debt to value ratio is given by Debt / Value of the firm = $92,000 / ( $92,000 + $2575) = 0.973.

(d) If the growth rate is 6% then the Earnings next year will be EBIT * ( 1 + g), where g is the growth rate.

= $6,440 * ( 1 + 0.06) = $6286.40.

Now the payment to shareholders will be (Earnings next year - EBIT = $6286.40 - $6,440 = $386.40

Now the value of equity will be Payment to shareholders / ( Rate of return - Growth rate)

= $386.40 / ( 0.07 - 0.06) = $3864.

   Debt to value ratio is given by Debt / Value of the firm = $92,000 / ( $92,000 + $6286.40) = 0.936.

Note:- If this answer helps you pls give thumbs up.